# A Comprehensive Survey of Synthetic Tabular Data Generation

### v2

@@ V-B3 Alignment
Compared to other evaluation matrices, alignment with background knowledge has recently attracted growing attention from the research community and is increasingly recognized as a critical criterion for evaluating the authenticity of synthetic data. There are three primary metrics commonly used to evaluate alignment in synthetic tabular data: (1) Constraint Violation Rate (CVR), which measures the percentage of synthetic samples that violate at least one predefined constraint; (2) Constraint Violation Coverage (CVC), which calculates the proportion of constraints that are violated by at least one sample in the dataset; and (3) Sample-wise Constraint Violation Coverage (sCVC), which computes the average percentage of samples that violate each individual constraint.
V-C Privacy.
Another important objective of synthetic tabular data is to enable data generation in privacy-sensitive domains, such as finance and healthcare, without disclosing sensitive information that could potentially identify individuals from the original dataset. Therefore, it is essential to evaluate the synthetic data using privacy metrics to ensure that privacy preservation is achieved. These metrics aim to quantify the extent to which the original records in the real dataset can be re-identified or inferred from the synthetic data. They are designed to assess whether the synthetic data inadvertently leaks sensitive information, either through direct resemblance to individual records or through indirect statistical inference. The primary privacy evaluation techniques include:
@@
  • Attribute Inference Attack: This type of privacy attack assumes that an adversary has partial knowledge of an individual’s data and attempts to infer missing or sensitive attributes using the synthetic tabular dataset $D_{s}$. The success of such an attack reflects the extent to which private information can be predicted from synthetic data. High inference accuracy indicates poor privacy preservation and potential leakage of attribute-level information.
  • Membership Inference Attack: In this setting, an attacker tries to determine whether a specific individual was part of the training data used to generate the synthetic tabular dataset $D_{s}$. This is typically done by exploiting differences in model behavior on seen versus unseen data. A high success rate in membership inference suggests that the generative model may have overfit the training data, leading to a privacy risk.
@@ VI-A Practical Applications
In this section, we examine the practical impact of synthetic tabular data generation from two distinct yet complementary perspectives: its use in improving tabular data availability and its application in privacy-preserving data analysis.
@@ VI-A1 Improving Tabular Data Availability
In practical applications, tabular data often suffers from class imbalance and missing values. These challenges can significantly impair the performance of downstream machine-learning tasks. As a result, improving data availability through the use of synthetic data has emerged as an essential strategy to alleviate such limitations and improve model robustness. For cases of imbalanced categories, models such as cGAN [48], CTAB-GAN [49], and CTGAN [38] incorporate conditional vectors into the GAN framework to guide the generation process and specifically synthesize samples from minority classes. SOS [65] proposes a style-transfer-based oversampling method for handling imbalanced tabular data using score-based generative models (SGMs). The core idea is to transform majority-class samples into minority-class samples through a score-based diffusion process, thereby addressing class imbalance in a more targeted and data-efficient manner. EPIC [71] and LITO [72] leverage prompt engineering to instruct LLMs to generate synthetic samples from minority classes. By crafting task-specific prompts, these methods effectively address class imbalance in tabular datasets through controlled, LLM-driven data generation. For the case of missing values, one common approach for handling missing values is Multiple Imputation by Chained Equations (MICE), an iterative method that estimates the conditional distribution of each feature given the others and imputes missing values accordingly. LLMs that have been fine-tuned on tabular data possess an inherent ability to infer missing values by leveraging prompt-based generation, without the need for retraining or explicit imputation models.
@@ VI-A2 Privacy Protection
Additionally, in domains such as finance and healthcare, strict compliance with data privacy regulations is required to ensure the security and confidentiality of user data. As a result, the use of synthetic data for privacy preservation has become particularly important in these sensitive application areas. medGAN and medBGAN apply generative adversarial networks (GANs) to the synthesis of electronic health records (EHRs). The effectiveness of the generated synthetic tabular data has been demonstrated through a combination of statistical distribution comparisons, predictive modeling tasks, and expert evaluations by medical professionals. These studies also show that synthetic data poses limited risks in terms of identity and attribute disclosure, thereby supporting its utility in privacy-sensitive healthcare applications. EHRDiff, TabDDPM-EHR, MedDiff, and DPM-EHR explore the use of diffusion models for synthetic tabular data generation in order to produce more realistic and high-fidelity synthetic electronic health records (EHRs). These methods leverage the denoising capabilities and distributional flexibility of diffusion-based generative frameworks to better capture the complex patterns inherent in medical tabular data. FLEXGEN-EHR further advances this line of research by explicitly addressing the challenge of missing modalities in EHRs. It introduces a unified learning framework that aligns heterogeneous data representations and effectively imputes missing information, thereby enhancing the completeness and utility of the generated synthetic records.
@@ VI-B Future Directions
  • Trade-off Between Data Availability and Privacy Protection. Many early methods have prioritized either data utility or privacy preservation, often resulting in highly specialized models with limited generalizability. This highlights a key challenge in the field: achieving a balanced trade-off between data utility and privacy protection.

### v3

@@ V-B3 Alignment
Alignment with background knowledge has recently attracted growing attention from the research community and is increasingly recognized as a critical criterion for evaluating the authenticity of synthetic data. There are three primary metrics commonly used to evaluate alignment in synthetic tabular data: (1) Constraint Violation Rate (CVR), which measures the percentage of synthetic samples that violate at least one predefined constraint; (2) Constraint Violation Coverage (CVC), which calculates the proportion of constraints that are violated by at least one sample in the dataset; and (3) Sample-wise Constraint Violation Coverage (sCVC), which computes the average percentage of samples that violate each individual constraint.
TABLE V: The performance of various methods for synthesizing tabular data is evaluated using multiple metrics. AUC and RMSE are used to assess the ML Efficiency of the synthesized data on classification and regression tasks, respectively. CVR evaluates how well the synthetic data aligns with background knowledge. Column-wise and Pair-wise metrics are employed to measure the fidelity of the synthesized data. The best results are highlighted in bold.
Data Metrics Real Method
CTGAN CTGAN+ TVAE P-TA TabDDPM TABSYN
[…]
V-C Privacy Protection
Another important objective of synthetic tabular data is to enable data generation in privacy-sensitive domains, such as finance and healthcare, without disclosing sensitive information that could potentially identify individuals from the original dataset. Therefore, it is essential to evaluate the synthetic data using privacy protection metrics to ensure that privacy protection is achieved. These metrics aim to quantify the extent to which the records in the real dataset can be re-identified or inferred from the synthetic data. They are designed to assess whether the synthetic data inadvertently leaks sensitive information, either through direct resemblance to individual records or through indirect statistical inference. The primary privacy protection evaluation techniques include:
@@ V-C Privacy Protection
  • Attribute Inference Attack: This type of privacy attack assumes that an adversary has partial knowledge of an individual’s data and attempts to infer missing or sensitive attributes using the synthetic tabular dataset $D_{s}$. The success of such an attack reflects the extent to which private information can be predicted from synthetic data. High inference accuracy indicates poor privacy protection and potential leakage of attribute-level information.
  • Membership Inference Attack: In this setting, an attacker tries to determine whether a specific individual was part of the training data used to generate the synthetic tabular dataset $D_{s}$. This is typically done by exploiting differences in model behavior on seen versus unseen data. A high success rate in membership inference suggests that the generative model may have overfit the training data, leading to a privacy protection risk.
@@
V-D Analysis and Discussion
To intuitively demonstrate the performance of different synthetic tabular data generation methods, we select a representative model from each summarized category and conduct experiments on several classic datasets. The experimental results are then analyzed and discussed. The results are shown in Table V. Compared to traditional generation methods, tabular data produced by Diffusion Model methods and LLM-based approaches exhibits markedly higher data availability. Nonetheless, the hallucination problem associated with LLMs poses a challenge to fidelity, especially in terms of pairwise fidelity. Diffusion models exhibit superior performance in generating synthetic tabular data with a small number of columns, while LLM-based methods are more effective in handling tabular data with a greater number of columns. While the diffusion model excels in processing numerical features, its performance advantage diminishes on datasets characterized by a higher proportion of categorical features, such as News and Shoppers.
@@ VI-A Practical Applications
In this section, we examine the practical impact of synthetic tabular data generation from two distinct yet complementary perspectives: its use in improving tabular data availability and its application in privacy protection data analysis.
@@ VI-A1 Improving Tabular Data Availability
In practical applications, tabular data often suffers from class imbalance and missing values. These challenges can significantly impair the performance of downstream machine-learning tasks. As a result, improving data availability through the use of synthetic data has emerged as an essential strategy to alleviate such limitations and improve model robustness. For cases of imbalanced categories, models such as cGAN [53], CTAB-GAN [54], and CTGAN [43] incorporate conditional vectors into the GAN framework to guide the generation process and specifically synthesize samples from minority classes. SOS [70] proposes a style-transfer-based oversampling method for handling imbalanced tabular data using score-based generative models (SGMs). The core idea is to transform the majority-class samples into minority-class samples through a score-based diffusion process, thereby addressing class imbalance in a more targeted and data-efficient manner. EPIC [76] and LITO [77] leverage prompt engineering to instruct LLMs to generate synthetic samples from minority classes. By crafting task-specific prompts, these methods effectively address class imbalance in tabular datasets through controlled, LLM-driven data generation. For the case of missing values, one common approach for handling missing values is Multiple Imputation by Chained Equations (MICE) [123], an iterative method that estimates the conditional distribution of each feature given the others and imputes missing values accordingly. Diffusion Model Methods [15, 61] typically generates missing values by conditioning on the column names of the missing entries. LLMs that have been fine-tuned on tabular data possess an inherent ability to infer missing values by leveraging prompt-based generation, without the need for retraining or explicit imputation models [76, 16, 79].
@@ VI-A2 Privacy Protection
Additionally, in domains such as finance and healthcare, strict compliance with data privacy protection regulations is required to ensure the security and confidentiality of user data. As a result, the use of synthetic data for privacy protection has become particularly important in these sensitive application areas. medGAN [45] and medBGAN [124] apply generative adversarial networks (GANs) to the synthesis of electronic health records (EHRs). The effectiveness of the generated synthetic tabular data has been demonstrated through a combination of statistical distribution comparisons, predictive modeling tasks, and expert evaluations by medical professionals. These studies also show that synthetic data poses limited risks in terms of identity and attribute disclosure, thereby supporting its utility in privacy-sensitive healthcare applications. EHRDiff [125], TabDDPM-EHR [64], MedDiff [126], and DPM-EHR [66] explore the use of diffusion models for synthetic tabular data generation in order to produce more realistic and high-fidelity synthetic electronic health records (EHRs). These methods leverage the denoising capabilities and distributional flexibility of diffusion-based generative frameworks to better capture the complex patterns inherent in medical tabular data. FLEXGEN-EHR [68] further advances this line of research by explicitly addressing the challenge of missing modalities in EHRs. It introduces a unified learning framework that aligns heterogeneous data representations and effectively imputes missing information, thereby enhancing the completeness and utility of the generated synthetic records.
@@ VI-B Future Directions
  • Trade-off Between Data Availability and Privacy Protection. Many early methods have prioritized either data utility or privacy protection, often resulting in highly specialized models with limited generalizability. This highlights a key challenge in the field: achieving a balanced trade-off between data utility and privacy protection.